Considered the reaction:





3 NaOH(aq)+ FeCI3(s) →3 NaCI (aq) + Fe(OH)3 (s)





(molar mass: Fe(OH)3=106.84 g/mole; FeCI3=162.19 g/mol






1) if 300 ml of 0.150 M NaOH were added to a solution containing excess FeCI3, CALCULATE the mass of Fe(OH)3 would be produced

1
Expert's answer
2022-03-12T02:00:06-0500

Moles of NaOH used are

(300/1000)×0.15= 0.045moles

Mole fraction= 3:1

Mole of Fe(OH)3 produced are

(1/4)× 0.045= 0.01125moles

"Mass = moles \u00d7 RMM"

Mass= 0.01125× 106.84

= 1.2g